OpenStudy (anonymous):

Express the derivative as a function of x alone. Derivative at listed in comments.

OpenStudy (anonymous):

\[y'=\frac{ 6x^2-y-1 }{ x }\]

OpenStudy (loser66):

don't get what you mean

OpenStudy (anonymous):

It is a 3 part problem, c is "Confirm that the two results are consistent by expressing the derivative in part (a) as a function of x alone." For part (a), the derivative is the y prime i stated in the comment.

OpenStudy (anonymous):

Do you mean to take the partial derivative?

OpenStudy (anonymous):

Sorry, I don't exactly understand what you mean by expressing it as a function of x alone.

OpenStudy (loser66):

shoot the original problem , please

OpenStudy (anonymous):

Ok, \[1) x+xy-2x^3=2\] a)Find dy/dx by differentiating implicitly. b)Solve the equation for y as a function of x, and find dy/dx from that equation. c)Confirm that the two results are consistent by expressing the derivative in part (a) as a function of x alone.

OpenStudy (loser66):

a)? show me your work, please

OpenStudy (anonymous):

\[1+y+y'x-6x^2=0\] \[y'=\frac{ 6x^2-y-1 }{ x }\]

OpenStudy (loser66):

b?

OpenStudy (anonymous):

\[y=\frac{ -x+2x^3+2 }{ x }\]

OpenStudy (anonymous):

Ah, I see what you mean now. I'll let Loser66 help on this one, though.

OpenStudy (anonymous):

\[y'=\frac{ 4x^3-2 }{ x^2 }\]

OpenStudy (loser66):

@bloopman I am not good at explanation, you help her, please

OpenStudy (loser66):

@asapthe for a, you still have y in y' right? replace that y = ... from original function, then, simplify it , you will see both a, b are the same.

OpenStudy (loser66):

got me?

OpenStudy (anonymous):

You have it right, Loser66.

OpenStudy (anonymous):

what do you mean y in y'?

OpenStudy (loser66):